# TrustSQL: Benchmarking Text-to-SQL Reliability with Penalty-Based Scoring

### v3

@@ 1 Introduction
Text-to-SQL is a task that involves generating SQL queries from natural language (NL) questions, a challenge that has persisted for several decades shwartz2022tabular . This task is crucial both within and beyond the realm of machine learning and natural language processing (NLP), due to the ubiquity of databases in various fields such as medicine, finance, retail, and telecommunications. Accessing structured data stored in these databases through natural language can unlock the true potential of big data, enabling domain experts to uncover new quantitative insights van2024tabular .
To effectively utilize these models, users need to be familiar with their capabilities. These include the types of questions each individual model can accurately answer, the information it has access to, and what operations can or cannot be performed via SQL. However, many potential users who could greatly benefit from the model are unfamiliar with these capabilities and may find it challenging to use, as their questions may frequently fall outside the model’s scope (beyond the correct SQL in Figure 1(b)). The challenge intensifies in real-world deployment scenarios where user questions include not only those feasible for the model but also infeasible ones, such as requests requiring knowledge beyond the database content or SQL functionalities (as shown in Figure 1(a)). Moreover, the absence of abstention mechanisms in current text-to-SQL modeling can lead to incorrect model outputs for both feasible and infeasible questions going unnoticed, potentially outweighing the benefits of model use. If a model, by design, could selectively answer correct SQL queries while abstaining from answering the rest, it would significantly boost user confidence, thereby enhancing trust in the model and encouraging its use whitehead2022reliable ; ren2022out .
@@
To address the aforementioned challenges, we present TrustSQL, a new benchmark designed to assess text-to-SQL reliability. We define text-to-SQL reliability as a model’s ability to handle any input questions correctly through accurate SQL generation and abstention, thereby reducing the red and yellow regions by covering the entire region in green, as shown in Figure 1(c)). Unlike standard text-to-SQL tasks defined within the feasible region (Figure 1(b)), real user questions span both feasible and infeasible task regions, including requests beyond the database content and SQL functionalities (Figure 1(a)). We constructed this new comprehensive setting using three domain-specific datasets—ATIS atis , Advising advising , and EHRSQL lee2022ehrsql —along with one cross-domain dataset—Spider yu2018spider . In constructing TrustSQL, we reflected common error cases undermining text-to-SQL reliability through careful data splitting and incorporation of various infeasible questions, similar to red teaming. We then evaluated existing methods using a novel penalty-based scoring metric designed specifically to measure text-to-SQL reliability (details in Section 3.2). Our experimental results reveal that while most advanced methods seem to perform well in penalty-less settings, achieving high scores under severe penalty settings remains highly challenging. From the high penalty perspective, knowing when to abstain precisely carries more weight than rashly attempting to generate SQL queries, underscoring the need for substantial efforts towards gaining trust in model outputs.
@@ 2 Dataset Construction
  • Complexity of SQL generation: User questions, especially those posed by domain experts, are often too complex for models to answer accurately. This complexity arises from the inherent difficulty of the queries yu2018spider ; li2023can ; hazoom2021sede ; lee2022ehrsql , as in Table 6 in Supplementary B, or from the model encountering unfamiliar phrases and entities specific to the database suhr-etal-2020-exploring ; lee-etal-2021-kaggledbqa . Models that generate SQL without the ability to abstain from such cases risk compromising trust in their output. We include these dimensions of query difficulty [P1] and familiarity [P2] in feasible data construction.
@@
  • Schema incompatibility: Users who are unfamiliar with the exact database schema may inadvertently ask questions that cannot be processed because the information they seek does not exist in the existing columns. Previous works construct such cases by deleting and perturbing existing columns triagesql ; known_dte . Instead, we include three distinct types within this case [P3] to provide more granular views on model behavior.
@@ 2.1 Feasible Data Construction
To establish the benchmark data covering comprehensive settings, we selected three complex domain-specific text-to-SQL datasets, namely, ATIS (flight information) atis , Advising (education) advising , and EHRSQL (healthcare) (lee2022ehrsql, ). Additionally, we included one cross-domain dataset, Spider. The distinction between domain-specific and cross-domain text-to-SQL settings, along with the complexity of each dataset, is detailed in Supplementary B.
@@
Our preliminary data analysis revealed that ATIS and Advising require significant preprocessing due to issues such as numerous query duplications, inconsistencies, and poor-quality natural language (NL) questions (each issue discussed in SupplementaryG). We implemented several strategies to correct these issues, including creating new question paraphrases for all input NL questions using GPT-4, followed by a manual review by the authors. For further details, please refer to Supplementary G). The databases of the four datasets and their corresponding question-SQL pairs, after preprocessing, are included as questions in the feasible task region (light blue region in Figure 1(a)). To reflect the query difficulty [P1], we adopted the SQL hardness criteria following pourreza2023din-sql across all datasets. Detailed statistics for each hardness level are provided in Table 6 in Supplementary B. Query familiarity [P2] is later introduced in the data splitting process (§2.3).
@@ 2.2 Infeasible Data Construction
[P3]: Instead of randomly dropping existing columns in the database triagesql ; known_dte , we introduce the concept of hypothetical columns as target columns for infeasible questions. Those questions are asked as if their corresponding SQL queries exist, but in fact do not, since no such columns exist in the database. We categorized these into three types: column-surface, column-related, and column-unrelated, representing attacks based on surface-form and semantic column similarities and dissimilarities, respectively.
@@
  • column-related: Questions referring to hypothetical columns that could naturally integrate into the existing database schema (e.g., a hypothetical column named ‘duration’ in the ‘song’ table). This question type can be seen as a manual version of the ADD perturbations in ADVETA adveta .
@@ 4.1.1 Pipeline-based Approach
In conventional NLP task formulations, enhancing text-to-SQL reliability can be addressed by combining three different tasks: infeasible question detection triagesql ; known_dte , SQL generation pourreza2023din-sql ; wang2023mac , and SQL error detection chen2023teaching .
@@
For SQL generation, we use T5 (T5-3B), feeding questions followed by a serialized schema listing all tables and columns suhr-etal-2020-exploring . To incorporate abstention mechanisms, we use SQLCoder (sqlcoder-7b-2) for infeasible question detection (before SQL generation) and error detection (after SQL generation). For infeasible question detection, we fine-tune SQLCoder on TriageSQL triagesql , denoted as CLSM. For error detection, we train SQLCoder based on the generated SQLs and their sample-level accuracy from trained T5 on the TrustSQL validation set plus samples of infeasible questions in TriageSQL, denoted as ErrorM. Details about training these models are provided in Supplementary C. When denoting the combined methods, we use arrows to indicate data flow between models, such as CLSMT5 indicating that question detection comes first, followed by SQL generation.
@@
For SQL generation, we employ three methods using GPT-3.5 (gpt-3.5-turbo-0125): SQLPrompt, SQL generation with few-shot prompting chang2023prompt , DIN-SQL pourreza2023din-sql , a state-of-the-art task decomposition framework in Spider, and MAC-SQL wang2023mac , a state-of-the-art multi-agent framework in BIRD li2023can . We adjust text prompts for DIN-SQL and MAC-SQL for domain-specific datasets. For abstention, we use GPT-4o for both infeasible question detection and error detection. We provide the modeling details in Supplementary D.
@@ 4.1.2 Unified Approach
For SQL generation, we use T5 (T5-3B) due to its wide use in selective generation ren2022out and calibration stengel2023calibrated in NLP. For abstention, we leverage the hidden representations of the input and the output probabilities by adopting uncertainty estimation to detect samples that are out-of-distribution (OOD) for abstention. Below are the methods we use:
  • Entropy-based: Each sample is assigned the maximum entropy value among the generated SQL tokens, denoted as MaxEnt (lee2022ehrsql, ).
  • Probability-based: Each sample is assigned the minimum of the highest token probabilities from the token sequence, denoted as MaxProb (stengel2023calibrated, ).
  • Distance-based: Mahalanobis distance calculation based on the distribution of the training data representations, denoted as FeatMD (lee2018simple, ; ren2022out, ), and relative Mahalanobis distance calculation with a background distribution using the TriageSQL dataset, denoted as FeatRMD (ren2021simple, ; ren2022out, ).
@@
  • SQLPrompt[DEMO]: Few-shot demonstrations of similar feasible questions with their corresponding SQL queries and similar infeasible questions, each labeled with ‘not_answerable’ sun2021conditionalqa . The order of demonstrations is randomly chosen.
  • SQLPrompt[Voting]: Using the same prompts as SQLPrompt, we sample the answers five times (wang2022self, ) and consider the SQL generation only if all samples unanimously agree; otherwise, we abstain.
@@ 6 Related Works
Text-to-SQL has attracted significant attention for its practical applications. Key single-domain datasets include ATIS (atis, ), GeoQuery (geography, ), Advising advising , SEDE hazoom2021sede , and EHRSQL lee2022ehrsql . For cross-domain tasks, Spider yu2018spider and BIRD li2023can serve as essential benchmarks for SQL generation. However, these benchmarks assume that all input questions should generate SQL, lacking incentives for developing models to abstain. This absence of safety measures undermines reliable deployment. Some efforts focus on detecting infeasible questions given database schemas triagesql ; known_dte , but they do not integrate SQL generation or consider the error-handling capabilities of text-to-SQL models. Our work aims to integrate comprehensive reliability measures, aligning the benchmark more closely with model trust.
@@
Uncertainty estimation is vital for AI reliability hendrycks2021unsolved . The methods include probability-based methods (calculating entropies or model-assigned probabilities)marek2021oodgan ; xiao2021hallucination ; kim2022uncertainty ; stengel2023calibrated and distance-based methods (using metrics such as the Mahalanobis distance to compare new inputs with training data)lee2018simple ; zheng2020out ; ren2021simple ; ren2022out . Probability-based methods gauge model confidence, enhancing output reliability. Distance-based methods identify outliers likely to produce low-quality outputs, promoting safer AI deployment. Our work employs both approaches to develop abstention mechanisms for text-to-SQL models.

### v4

@@ 1 Introduction
Text-to-SQL is a task that involves generating SQL queries from natural language (NL) questions, a challenge that has persisted for several decades [17]. This task is crucial both within and beyond the realm of machine learning and natural language processing (NLP), due to the ubiquity of databases in various fields such as medicine, finance, retail, and telecommunications. Accessing structured data stored in these databases through natural language can unlock the true potential of big data, enabling domain experts to uncover new quantitative insights [21].
To effectively utilize these models, users need to be familiar with their capabilities. These include the types of questions each individual model can accurately answer, the information it has access to, and what operations can or cannot be performed via SQL. However, many potential users who could greatly benefit from the model are unfamiliar with these capabilities and may find it challenging to use, as their questions may frequently fall outside the model’s scope (beyond the correct SQL in Figure 1(b)). The challenge intensifies in real-world deployment scenarios where user questions include not only those feasible for the model but also infeasible ones, such as requests requiring knowledge beyond the database content or SQL functionalities (as shown in Figure 1(a)). Moreover, the absence of abstention mechanisms in current text-to-SQL modeling can lead to incorrect model outputs for both feasible and infeasible questions going unnoticed, potentially outweighing the benefits of model use. If a model, by design, could selectively answer correct SQL queries while abstaining from answering the rest, it would significantly boost user confidence, thereby enhancing trust in the model and encouraging its use [25, 16].
@@
To address the aforementioned challenges, we present TrustSQL, a new benchmark designed to assess text-to-SQL reliability. We define text-to-SQL reliability as a model’s ability to handle any input questions correctly through accurate SQL generation and abstention, thereby reducing the red and yellow regions by covering the entire region in green, as shown in Figure 1(c)). Unlike standard text-to-SQL tasks defined within the feasible region (Figure 1(b)), real user questions span both feasible and infeasible task regions, including requests beyond the database content and SQL functionalities (Figure 1(a)). We constructed this new comprehensive setting using three domain-specific datasets—ATIS [3], Advising [4], and EHRSQL [9]—along with one cross-domain dataset—Spider [27]. In constructing TrustSQL, we reflected common error cases undermining text-to-SQL reliability through careful data splitting and incorporation of various infeasible questions, similar to red teaming. We then evaluated existing methods using a novel penalty-based scoring metric designed specifically to measure text-to-SQL reliability (details in Section 3.2). Our experimental results reveal that while most advanced methods seem to perform well in penalty-less settings, achieving high scores under severe penalty settings remains highly challenging. From the high penalty perspective, knowing when to abstain precisely carries more weight than rashly attempting to generate SQL queries, underscoring the need for substantial efforts towards gaining trust in model outputs.
@@ 2 Dataset Construction
  • Complexity of SQL generation: User questions, especially those posed by domain experts, are often too complex for models to answer accurately. This complexity arises from the inherent difficulty of the queries [27, 11, 5, 9], as in Table 7 in Supplementary B, or from the model encountering unfamiliar phrases and entities specific to the database [19, 8]. Models that generate SQL without the ability to abstain from such cases risk compromising trust in their output. We include these dimensions of query difficulty [P1] and familiarity [P2] in feasible data construction.
@@
  • Schema incompatibility: Users who are unfamiliar with the exact database schema may inadvertently ask questions that cannot be processed because the information they seek does not exist in the existing columns. Previous works construct such cases by deleting and perturbing existing columns [29, 22]. Instead, we include three distinct types within this case [P3] to provide more granular views on model behavior.
@@ 2.1 Feasible Data Construction
To establish the benchmark data covering comprehensive settings, we selected three complex domain-specific text-to-SQL datasets, namely, ATIS (flight information) [3], Advising (education) [4], and EHRSQL (healthcare) [9]. Additionally, we included one cross-domain dataset, Spider. The distinction between domain-specific and cross-domain text-to-SQL settings, along with the complexity of each dataset, is detailed in Supplementary B.
@@
Our preliminary data analysis revealed that ATIS and Advising require significant preprocessing due to issues such as numerous query duplications, inconsistencies, and poor-quality natural language (NL) questions (each issue discussed in Supplementary G). We implemented several strategies to correct these issues, including creating new question paraphrases for all input NL questions using GPT-4, followed by a manual review by the authors. For further details, please refer to Supplementary G). The databases of the four datasets and their corresponding question-SQL pairs, after preprocessing, are included as questions in the feasible task region (light blue region in Figure 1(a)). To reflect the query difficulty [P1], we adopted the SQL hardness criteria following [14] across all datasets. Detailed statistics for each hardness level are provided in Table 7 in Supplementary B. Query familiarity [P2] is later introduced in the data splitting process (§2.3).
@@ 2.2 Infeasible Data Construction
[P3]: Instead of randomly dropping existing columns in the database [29, 22], we introduce the concept of hypothetical columns as target columns for infeasible questions. Those questions are asked as if their corresponding SQL queries exist, but in fact do not, since no such columns exist in the database. We categorized these into three types: column-surface, column-related, and column-unrelated, representing attacks based on surface-form and semantic column similarities and dissimilarities, respectively.
@@
  • column-related: Questions referring to hypothetical columns that could naturally integrate into the existing database schema (e.g., a hypothetical column named ‘duration’ in the ‘song’ table). This question type can be seen as a manual version of the ADD perturbations in ADVETA [13].
@@ 4.1.1 Pipeline-based Approach
In conventional NLP task formulations, enhancing text-to-SQL reliability can be addressed by combining three different tasks: infeasible question detection [29, 22], SQL generation [14, 23], and SQL error detection [2].
@@
For SQL generation, we use T5 (T5-3B), feeding questions followed by a serialized schema listing all tables and columns [19]. To incorporate abstention mechanisms, we use SQLCoder (sqlcoder-7b-2) for infeasible question detection (before SQL generation) and error detection (after SQL generation). For infeasible question detection, we fine-tune SQLCoder on TriageSQL [29], denoted as CLSM. For error detection, we train SQLCoder based on the generated SQLs and their sample-level accuracy from trained T5 on the TrustSQL validation set plus samples of infeasible questions in TriageSQL, denoted as ErrorM. Details about training these models are provided in Supplementary C. When denoting the combined methods, we use arrows to indicate data flow between models, such as CLSMT5 indicating that question detection comes first, followed by SQL generation.
@@
For SQL generation, we employ three methods using GPT-3.5 (gpt-3.5-turbo-0125): SQLPrompt, SQL generation with few-shot prompting [1], DIN-SQL [14], a state-of-the-art task decomposition framework in Spider, and MAC-SQL [23], a state-of-the-art multi-agent framework in BIRD [11]. We adjust text prompts for DIN-SQL and MAC-SQL for domain-specific datasets. For abstention, we use GPT-4o for both infeasible question detection and error detection. We provide the modeling details in Supplementary D.
@@ 4.1.2 Unified Approach
For SQL generation, we use T5 (T5-3B) due to its wide use in selective generation [16] and calibration [18] in NLP. For abstention, we leverage the hidden representations of the input and the output probabilities by adopting uncertainty estimation to detect samples that are out-of-distribution (OOD) for abstention. Below are the methods we use:
  • Entropy-based: Each sample is assigned the maximum entropy value among the generated SQL tokens, denoted as MaxEnt [9].
  • Probability-based: Each sample is assigned the minimum of the highest token probabilities from the token sequence, denoted as MaxProb [18].
  • Distance-based: Mahalanobis distance calculation based on the distribution of the training data representations, denoted as FeatMD [10, 16], and relative Mahalanobis distance calculation with a background distribution using the TriageSQL dataset, denoted as FeatRMD [15, 16].
@@
  • SQLPrompt[DEMO]: Few-shot demonstrations of similar feasible questions with their corresponding SQL queries and similar infeasible questions, each labeled with ‘not_answerable’ [20]. The order of demonstrations is randomly chosen.
  • SQLPrompt[Voting]: Using the same prompts as SQLPrompt, we sample the answers five times [24] and consider the SQL generation only if all samples unanimously agree; otherwise, we abstain.
@@ 6 Related Works
Text-to-SQL has attracted significant attention for its practical applications. Key single-domain datasets include ATIS [3], GeoQuery [28], Advising [4], SEDE [5], and EHRSQL [9]. For cross-domain tasks, Spider [27] and BIRD [11] serve as essential benchmarks for SQL generation. However, these benchmarks assume that all input questions should generate SQL, lacking incentives for developing models to abstain. This absence of safety measures undermines reliable deployment. Some efforts focus on detecting infeasible questions given database schemas [29, 22], but they do not integrate SQL generation or consider the error-handling capabilities of text-to-SQL models. Our work aims to integrate comprehensive reliability measures, aligning the benchmark more closely with model trust.
@@
Uncertainty estimation is vital for AI reliability [6]. The methods include probability-based methods (calculating entropies or model-assigned probabilities)[12, 26, 7, 18] and distance-based methods (using metrics such as the Mahalanobis distance to compare new inputs with training data)[10, 30, 15, 16]. Probability-based methods gauge model confidence, enhancing output reliability. Distance-based methods identify outliers likely to produce low-quality outputs, promoting safer AI deployment. Our work employs both approaches to develop abstention mechanisms for text-to-SQL models.